Differentiate the following functions (i) 𝑓(π‘₯) = π‘₯𝑒 π‘₯ 𝑐𝑠𝑐π‘₯, (ii) β„Ž(π‘₯) = ln(π‘₯ + √π‘₯ 2 βˆ’ 1), (iii) 𝑦 = cos(𝑒 βˆšπ‘‘π‘Žπ‘› 3π‘₯ ), (iv) √1 + sin3(π‘₯𝑦 2) = 𝑦, (v) 𝑦(π‘₯) = (log𝑒 π‘₯) π‘₯ tan(𝑒 π‘₯ ), (vi) π‘š(π‘₯) = sinβˆ’1 (√1 βˆ’ 9π‘₯ 2).Β 


1
Expert's answer
2021-09-28T00:11:43-0400

i)

"f(x)=xe^xcscx"

"f'(x)=e^x(cscx-xcosx\/sin^2x)+xe^xcscx"


ii)

"h(x)=ln(x+\sqrt{x^2-1})"


"h'(x)=\frac{1+x/\sqrt{x^2-1}}{x+\sqrt{x^2-1}}"


iii)

"y=cos(e^{\sqrt{tan(3x)}})"


"y'=-3sin(e^{\sqrt{tan(3x)}})sec^2(3x)e^{\sqrt{tan(3x)}}/(2\sqrt{tan(3x)})"


iv)

"y=\sqrt{1+sin^3(xy^2)}"


"y'=\frac{3sin^2(xy^2)cos(xy^2)(y^2+2xyy')}{2\sqrt{1+sin^3(xy^2)}}"


"y'=\frac{3sin^2(xy^2)cos(xy^2)y^2}{2\sqrt{1+sin^3(xy^2)}}/(1-\frac{3sin^2(xy^2)cos(xy^2)(y^2+2xy)}{2\sqrt{1+sin^3(xy^2)}})"


v)

"y(x)=xlog_extan(e^x)=xlnxtan(e^x)"

"y'(x)=tan(e^x)(1+lnx)+xe^xlnxsec^2(e^x)"


vi)

"m(x)=sin^{-1}(\sqrt{1-9x^2})"


"m'(x)=-\frac{18x}{6x\sqrt{1-9x^2}}=-\frac{3}{\sqrt{1-9x^2}}"